# The Preheating Stage on The Starobinsky Inflation after ACT

### v3

@@ Abstract
In this paper, we reinvestigate the Starobinsky inflation model and its reheating features in light of the recent ACT results. To make the Starobinsky model consistent with the ACT data at the $68\%$ confidence level, the number of e-folds must increase while the reheating temperature decreases. We find that the Starobinsky model requires a spectator field to achieve efficient preheating. The preheating stage and the reheating temperature must be significantly adjusted to accommodate the lower temperature. In this paper, the favored non-minimal coupling of the produced particles is approximately $10$ or slightly lower. We also present viable parameter sets that fit the preferred reheating mechanism in this model. For certain parameter choices, the daughter fields could potentially be detected in future collider experiments such as the LHC or the ILC. Furthermore, our proposed mechanism can reproduce the lower reheating temperature, but it fails when the temperature falls below $1$ GeV.
@@ 1 Introduction
The Starobinsky model [starobinsky], as the simplest example of a modified gravity theory [de2010f, sotiriou2010f, nojiri2011unified], is considered one of the most successful inflationary models. Its simplicity lies in the fact that it relies solely on gravity as the driving mechanism for inflation. Despite this simplicity, it remains consistent with the Cosmic Microwave Background (CMB) observations from Planck data [cmb]. In recent years, this model and its various extensions have continued to attract considerable attention and discussion [kehagias, Toyama:2024ugg, Ketov:2024klm, Ketov:2025nkr, He:2024wqv, He:2018gyf, Risdianto:2022ium, Sebastiani:2013eqa, Bamba:2014jia, Nojiri:2017ncd, CruzDombriz:2016wcb, Odintsov:2018zhw, Odintsov:2022hxu, Odintsov:2021kup, Odintsov:2020thl, Nojiri:2019fft, Elizalde:2018rmz]. However, the Starobinsky model is now challenged by newly established observational data from the Atacama Cosmology Telescope (ACT) [louis2025atacama]. This new data shows that the slow-roll parameter is $n_{s}=0.9709\pm 0.0038$. When combined with measurements from CMB lensing, Baryon Acoustic Oscillation (BAO), and the Dark Energy Spectroscopic Instrument (DESI), the extended result yields $n_{s}=0.9743\pm 0.0034$. This result indicates that the Starobinsky model is excluded at the $2\sigma$ level for $N\approx 60$. The model remains viable if a large number of e-folds ($N\approx 75$) is allowed, which leads to a lower reheating temperature [Drees:2025ngb, liu2025reconciling, mohammadi2025starobinsky, zharov2025reheating, gialamas2025keeping]. Alternatively, this issue can be addressed by introducing higher-curvature perturbations to satisfy the new constraint [addazi2025curvature].
@@
The reheating temperature of the Starobinsky model is found to be relatively low $\sim 10^{9}$ GeV [de2010f, dorsch2024gravitational] compared to other models, such as Higgs inflation [bezrukov2009initial]. This lower reheating temperature is often preferred in inflationary cosmology, as it helps avoid the overproduction of unwanted relics [kallosh2000gravitino]. For example, Refs. [ellis1982inflation, kawasaki1995gravitino] argue [^1] See also Ref. [mcdonald2000reheating]. that the reheating temperature should be below $10^{9}$ GeV to prevent excessive production of gravitino. Such a low reheating temperature significantly influences the thermal history of the universe, particularly with respect to Baryogenesis [davidson2000baryogenesis] and the production of dark matter. However, as for leptogenesis, it requires at least several MeV of reheating temperatures to operate successfully [hannestad2004lowest, choi2017new]. In summary, estimation of the reheating temperature range is wide, from a few MeV up to nearly the Planck scale. This broad range suggests that the reheating temperature is still poorly constrained. In addition to the ACT observation results, the Starobinsky inflation model is expected to yield a much lower reheating temperature.
@@
The preheating stage of the Starobinsky model has received comparatively less attention in the literature than its inflationary phase. This is largely because the inflationary dynamics are better constrained, whereas the preheating phase tends to be more model-dependent [van2017reheating, fu2019nonlinear]. [^2] One can see the historical preheating stage in Refs. [Dolgov:1989us, Traschen:1990sw, Shtanov:1994ce, Kofman:1994rk] In the Starobinsky model, preheating is primarily driven by the interaction of the produced particles with gravity. While this makes it a simple and potentially promising scenario, our limited understanding of gravity makes it more challenging to analyze fully. In this paper, we investigate the preheating mechanism in the Starobinsky model, focusing on the efficient energy drain by preheating, the constraints from primordial black hole (PBH) production, and the resulting reheating temperature.
@@
This paper is organized as follows: In Section 2, we discuss the inflationary description of the Starobinsky model. In Section 3, we estimate the resulting reheating temperature by the ACT constraint. In Section 4, we examine the preheating phase of the Starobinsky model, which is non-minimally coupled to scalars. In Section 5, we consider the possibility of fermionic preheating in our model. We also discuss primordial black hole (PBH) formation in Section 6. The reheating scenario is discussed in Section 7. Finally, we present our conclusions in Section 8.
@@ 2 The inflationary model
$S_{E}=\int d^{4}x\sqrt{-g_{E}}\left[\frac{1}{2}M_{p}^{2}R_{E}-\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)\right],$ (2.3)
@@
$n_{s}=1-6\epsilon+2\eta,\quad r=16\epsilon,\quad A_{s}=\frac{V(\phi)}{24\pi^{2}\epsilon}\quad\epsilon=\frac{1}{2}M_{p}^{2}\left(\frac{V^{\prime}}{V}\right)^{2}\quad\eta=M_{p}^{2}\frac{V^{\prime\prime}}{V}.$ (2.5)
The prime denotes the derivative with respect to scalaron $\phi$. By using the ACT from the extended dataset P-ACT-LB yields [louis2025atacama] $n_{s}=0.9743\pm 0.0038$, $A_{s}\approx 2.13\times 10^{-9}$ at the pivot scale $\kappa_{*}=0.05$ Mpc-1. It can be used to determine the new constraint on the $\phi_{\text{ini}}$, which corresponds to the field value of the scalaron at $\kappa_{*}$, to be
@@
$N=\frac{1}{M_{p}^{2}}\int^{\phi_{\text{ini}}}_{\phi_{\text{end}}}\frac{V}{V^{\prime}}\simeq 75.51,$ (2.8)
much higher than the previously predicted  [^4] One should also note that slight changes in the constraint of $n_{s}$ could potentially alter the results of other parameters greatly. and this large number of e-folds is expected. [^5] The chosen e-fold greater than 60 is rather implausible but still acceptable [liddle2003long]. Note that we only used the center value of $n_{s}$. By using the range of $n_{s}$, we will obtain an e-fold range of 668766-87, which will later significantly alter the reheating temperature. However, to simplify our analysis, we will only consider the center value of $n_{s}$.This assumption is made to incorporate the new mechanism consistent with the lower reheating temperature indicated by the recent ACT results.
In addition, the effective mass of the scalaron $M$ is also affected by the new ACT constraint on the scalar power spectrum $A_{s}$ as
@@
which is slightly lower than if constrained by the Planck result [cmb].
@@ 3 The reheating constraint by ACT results
In this, we will estimate the reheating temperature with the new ACT results. This constrained reheating temperature will be important to estimate the proper reheating process as depicted in section 3. Straightforwardly, it is necessary to write the formula as
@@
where $a_{0}=1$ and $H_{0}=66.78\hskip 2.84526pt\text{km Mpc}^{-1}\text{s}^{-1}$ [louis2025atacama] represent the scale factor and Hubble parameter today. $a_{\kappa}$ and $H_{\kappa}$ represent the scale factor and Hubble parameter at the pivot scale $\kappa_{*}=0.05\hskip 2.84526pt\text{Mpc}^{-1}$. The subscripts ’eq’, ’end’, and ’reh’ correspond to ’matter-radiation equality’, ’end of inflation’, and ’reheating’, respectively. Note that we solve Eq. (3.1) by using $T_{\text{eq}}=0.8$ eV, and the end of preheating is evaluated at $\tilde{\phi}=M$. Finally, by using (3.1) with e-fold $N=75.5$, we obtain an upper bound for the reheating temperature as $T_{\text{reh}}\sim 10^{4}$ GeV. This value is five orders of magnitude lower than the previous prediction [de2010f]. The reheating temperature in our calculation is evaluated at the early radiation-dominated regime.
Note that we used the center value of $n_{s}$. This is to simplify our preheating mechanism. The small change in $n_{s}$ will greatly change the e-fold and finally alter the reheating temperature more severely. The reheating temperature should be directly connected to the preheating stage mechanisms. In that case,
@@ 4 The preheating
When $\phi\lesssim M_{P}$, the inflaton oscillates, marking the end of inflation. The particle produced by this mode is non-perturbative. If we assume $\chi$ is dominantly produced during this mode, which is presented by the action:
$S_{\chi}=\int d^{4}x\sqrt{-g_{J}}\left[\frac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\frac{1}{2}m_{\chi}^{2}\chi^{2}-\xi R_{J}\chi^{2}\right],$ (4.1)
we can evaluate its particle production. Firstly, we will calculate the analytical solution of $\phi$ as [kofman1997towards]
@@
$\ddot{\chi}_{k}+3H\dot{\chi}_{k}-3\sqrt{\frac{2}{3}}\frac{\dot{\phi}\dot{\chi}_{k}}{M_{p}}+\left[\frac{k^{2}_{\chi}}{a^{2}}+e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}m_{\chi}^{2}-3\xi M^{2}e^{2\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\left(e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}-1\right)\right]\chi_{k}=0,$ (4.3)
@@
on calculating (4.3). In Fig. 2, we show that the resonance production of $\chi_{k}$ is highly sensitive to both the non-minimal coupling $\xi$ and the initial condition $\chi_{k}(0)$. Our results indicate that particle production from zero-crossings alone is inefficient. Instead, efficient preheating relies primarily on tachyonic instability immediately after the first crossing. As we stated earlier, tachyonic preheating in the Starobinsky model is also strongly dependent on the initial conditions set at the start of preheating. If these conditions are not satisfied, Starobinsky inflation faces two possible results: (1) inconsistency with the observed particle abundance today (especially if $\chi$ is related to dark matter) or (2) prolonged preheating to generate the required number of particles.
@@
At the early stage of preheating, the inflaton perturbative decay and the scattering process are possible. The single perturbative decay, in principle, is more favored than scattering, as it needs a large density of incoherent inflaton particles to scatter. In contrast, at the start of preheating, the inflaton is mostly a coherent oscillating field. The perturbative decay itself is slow to be produced at the early stage of preheating. Additionally, if the decay occurs during inflation, it may spoil the slow-roll condition or leave observational relics. On the other hand, the favored $\chi\sim 10^{-2.5}M_{p}$ could be obtained from the fluctuation of $\delta\chi$. But it can only provide about $H/2\pi\sim 10^{-6}M_{p}$, in which about $10^{-3}$ of smaller order from the desired value. Lastly, the only possible way came from the fact that $\chi$ should be the spectator field during inflation. On the other hand, as $\chi$ is small-coupled with scalaron, it can be assumed it is safe as a spectator field. We can also conclude that for the successful preheating of the Starobinsky inflation, the appearance of the spectator field is necessary. At the early stage of preheating, in order to fulfill the non-relativistic conditions and the efficient preheating, $k_{\chi}$ should be extremely small. Thus, it gives $k_{\chi}^{2}\ll\xi M^{2}$.
Before proceeding, we clarify why the analytical time evolution of $\phi$ is used, even though it could in principle be numerically coupled with $\chi$ to obtain exact results. First, in the pure Starobinsky model, the scalaron $\phi$ is only weakly coupled to other fields during the initial stage of preheating. In the Einstein frame, the Lagrangian of the $\chi$ field is heavily suppressed by the factor $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}$, which indicates that $\phi$ and $\chi$ are weakly coupled. This justifies treating $\chi$ as a spectator field during inflation, so that the analytical form of Eq. (4.2) remains valid until $\phi$ becomes small. Also, we restrict our attention to the early phase, where $\phi$ is still not too small after the first crossing. In this regime, the analytical estimation of the time evolution of $\phi$ remains reliable. Second, solving the coupled equations of motion poses a significant challenge, as it involves complicated calculations, including integro-differential equations in $\chi^{2}$. Third, since our paper primarily focuses on the reheating mechanism in light of the new ACT results, we approximate preheating by considering how efficiently it can drain the inflaton field energy density. A more detailed analysis of the preheating stage may be presented in future work.
@@
The approximation given in Eq. (4.5) can be considered to be oversimplified. As shown in the Fig. 3, the analytical results are heavily deviated from Eq. (4.3). In that figure, we used the comparison of $\chi_{k}^{2}$ instead of $\chi_{k}$. In Fig. 3, we used the amplification factor of $\chi_{k}$ as $\sim 10$ (or $\sim 10^{2}$ for $\chi^{2}_{k}$) on the numerical result to be aligned with the analytical result. We will use this information to estimate the particle production during the crossing and the deviation factor of the Mathieu instability.
Figure 3: The comparison of analytical result (solid-red) and numerical result (solid-black) in $\chi_{k}^{2}$. We used amplification $1.1\times 10^{1.8}$ on the numerical result to match the analytical ones. In this plot, we used $\xi=10$, $k_{\chi}=10^{-8}M_{p}$, $m_{\chi}=10^{-8}M_{p}$, $\chi_{k}(0)=10^{2.5}M_{p}$ and $\dot{\chi}_{k}=0$.
For the near-zero crossing ($\sin(Mt)\simeq Mt$), we can approximate Eq. (4.5) as
@@
$\begin{split}&p^{2}=\left(k_{\chi}^{2}+m_{\chi}^{2}\right)\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{-2/3}\\ &q=\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{1/3}t.\end{split}$ (4.7)
@@
For $p^{2}\ll q$, we can estimate the particle production of $\chi$ by the first crossing of $\phi$ as [bezrukov2009initial, risdianto2025second]
@@
In this calculation, for the crossing to achieve the same energy as the inflaton ($\tfrac{1}{2}M^{2}\tilde{\phi}^{2}\sim 10^{-11}M_{p}^{4}$), we require $\xi\gg\mathcal{O}(1)$, which is strongly disfavored due to naturalness issues. In addition, Fig. 3 shows that the actual value of $\delta\rho_{\text{cross}}$ obtained analytically is about $\sim 10^{2}$ times smaller, indicating that the inflaton’s energy drain through zero-crossing becomes less efficient. It should also be noted that $\delta\rho_{\text{cross}}\propto|\chi_{k}|^{2}$.
@@
So we obtain the famous Mathieu equation as
@@
where $\theta_{\mu}$ is a phase. It is shown in the Mathieu instability band of Fig. 4 that certain parameters could potentially fall in the instability region in which the particle is largely produced. However, this semi-analytical result of the instability band could not show the effect of the initial conditions. Also, the parameter used in the numerical results could fit in the Fig. 4 with much smaller amplification by an order of 10.

### v4

@@ Abstract
In this paper, we reinvestigate the Starobinsky inflation model and its reheating features in light of the recent ACT results. To make the Starobinsky model consistent with the ACT data at the $68\%$ confidence level, the number of e-folds must increase while the reheating temperature decreases. We find that the Starobinsky model requires a spectator field to achieve efficient preheating. The preheating stage and the reheating temperature must be significantly adjusted to accommodate the lower temperature. In this paper, the favored non-minimal coupling of the produced particles is approximately $10$ or lower to avoid the unitarity issue. We also present viable parameter sets that fit the preferred reheating mechanism in this model. For certain parameter choices, the daughter fields could potentially be detected in future collider experiments such as the LHC or the ILC. Furthermore, our proposed mechanism can reproduce the lower reheating temperature, but it fails when the temperature falls below $1$ GeV.
@@ 1 Introduction
The Starobinsky model [59], as the simplest example of a modified gravity [9, 58, 46], is considered one of the most successful inflationary models. Its simplicity lies in the fact that it relies solely on gravity as the driving mechanism for inflation. Despite this simplicity, it remains consistent with the Cosmic Microwave Background (CMB) observations from PLANCK data [2]. In recent years, this model and its various extensions have continued to attract considerable attention and discussion [31, 60, 33, 32, 26, 27, 53, 56, 4, 44, 10, 50, 49, 48, 47, 45, 14]. However, the Starobinsky model is now challenged by newly established observational data from the Atacama Cosmology Telescope (ACT) [40]. This new data shows that the slow-roll parameter is $n_{s}=0.9709\pm 0.0038$. When combined with measurements from CMB lensing, Baryon Acoustic Oscillation (BAO), and the Dark Energy Spectroscopic Instrument (DESI), the extended result yields $n_{s}=0.9743\pm 0.0034$. This result indicates that the Starobinsky model is excluded at the $2\sigma$ level for $N\approx 60$. The model remains viable if a large number of e-folds ($N\approx 75$) is allowed, which leads to a lower reheating temperature [13, 39, 42, 63, 17]. Alternatively, this issue can be addressed by introducing higher-curvature perturbations to satisfy the new constraint [1].
@@
The reheating temperature of the Starobinsky model is found to be relatively low $\sim 10^{9}$ GeV [9, 12] compared to other models, such as Higgs inflation [6]. This lower reheating temperature is often preferred in inflationary cosmology, as it helps avoid the overproduction of unwanted relics [29]. For example, Refs. [15, 30] argue [^1] See also Ref. [41]. that the reheating temperature should be below $10^{9}$ GeV to prevent excessive production of gravitino. Such a low reheating temperature significantly influences the thermal history of the universe, particularly with respect to baryogenesis [8] and the production of dark matter. However, as for leptogenesis, it requires at least several MeV of reheating temperatures to operate successfully [24, 7]. In summary, estimation of the reheating temperature range is wide, from a few MeV up to nearly the Planck scale. This broad range suggests that the reheating temperature is still poorly constrained. In addition to the ACT observation results, the Starobinsky inflation model is expected to yield a much lower reheating temperature.
@@
The preheating stage of the Starobinsky model has received comparatively less attention in the literature than its inflationary stage. The explanation should be started as the fact that inflationary dynamics are better constrained, whereas the preheating phase tends to be strongly model-dependent [62, 16]. [^2] One can see the historical preheating stage in Refs. [11, 61, 57, 34] In the Starobinsky model, preheating is primarily driven by the interaction of the produced particles with gravity. While this scenario is simple and potentially promising, our limited understanding of gravity makes it more challenging to be favored as a realistic theory. In this paper, we investigate the preheating mechanism in the Starobinsky model, focusing on the efficient energy drain through preheating, the constraints imposed by primordial black hole (PBH) production, and the resulting reheating mechanism.
@@
This paper is organized as follows: In Section 2, we discuss the inflationary description of the Starobinsky model. In Section 3, we estimate the resulting reheating temperature by the ACT constraint. In Section 4, we examine the preheating phase of the Starobinsky model, which is non-minimally coupled to scalars. In Section 5, we consider the possibility of fermionic preheating in our model. We also discuss primordial black hole (PBH) formation in Section 6. The realistic reheating scenario is discussed in Section 7. Finally, we present our conclusions in Section 8.
@@ 2 The inflationary model
$S_{E}=\int d^{4}x\sqrt{-g_{E}}\left[\frac{1}{2}M_{p}^{2}R_{E}-\frac{1}{2}\partial^{\mu}\phi\partial_{\mu}\phi-V(\phi)\right],$ (2.3)
@@
$n_{s}=1-6\epsilon+2\eta,\quad r=16\epsilon,\quad\epsilon=\frac{1}{2}M_{p}^{2}\left(\frac{V^{\prime}}{V}\right)^{2},\quad\eta=M_{p}^{2}\frac{V^{\prime\prime}}{V},\quad A_{s}=\frac{V(\phi)}{24\pi^{2}\epsilon}.$ (2.5)
The prime denotes the derivative with respect to scalaron $\phi$. By using the ACT from the extended dataset P-ACT-LB yields [40] $n_{s}=0.9743\pm 0.0038$, $A_{s}\approx 2.13\times 10^{-9}$ at the pivot scale $\kappa_{*}=0.05$ Mpc-1. The $n_{s}$ can be used to determine the new constraint on the $\phi_{\text{ini}}$, which corresponds to the field value of the scalaron at $\kappa_{*}$, to be
@@
$N=\frac{1}{M_{p}^{2}}\int^{\phi_{\text{ini}}}_{\phi_{\text{end}}}\frac{V}{V^{\prime}}d\phi\simeq 75.51,$ (2.8)
much higher than the previously predicted by PLANCK results. It means, the increase on $n_{s}$ slightly by ACT results could increase the e-fold greatly. [^4] The choice of an e-fold number greater than 60 is rather implausible, although it remains acceptable [36]. Furthermore, the previous PLANCK results indicate that the Starobinsky model is favored by observations for approximately $60$ e-folds. In contrast, results from ACT favor significantly higher values of the number of e-folds (See results on Eq. (2.8)). Note that we only used the center value of $n_{s}=0.9743$ and we obtain e-fold $N=75.5$ as shown by the calculation from eq. (2.5)-(2.8). In addition, by using the range of $n_{s}$, we will obtain an e-fold range of 668766-87, which will later significantly alter the reheating temperature greatly. However, to simplify our analysis, we will only consider the center value of $n_{s}$. This assumption is made to incorporate the new mechanism consistent with the lower reheating temperature indicated by the recent ACT results. On the other hand, we will also discuss the possible range of reheating temperature if we keep the upper and lower bound of $n_{s}$. Description about the large e-fold and reheating temperature may clear in the section 3. The effective mass of the scalaron $M$ is also affected by the new ACT constraint on the scalar power spectrum $A_{s}$ as
@@
which is slightly lower than if constrained by the Planck result [2].
@@ 3 The reheating constraint by ACT results
In this paper, we will estimate the reheating temperature with the new ACT results. This constrained reheating temperature will be important to estimate the proper reheating process as depicted in section 7. Straightforwardly, it is necessary to write the formula as
@@
or in the more convenient form
$N=-\ln\left(\frac{\kappa_{*}}{a_{0}H_{0}}\right)+\frac{2}{3}\ln\left(\frac{M}{M_{p}}\right)+\ln\left(\frac{T_{\text{eq}}}{T_{\text{reh}}}\right)+\frac{1}{2}\ln\left(1+z_{\text{eq}}\right)+\frac{1}{2}\ln\left(\frac{\rho_{\kappa}}{\rho_{\text{eq}}}\right),$ (3.2)
where $a_{0}=1$ and $H_{0}=66.78\hskip 2.84526pt\text{km Mpc}^{-1}\text{s}^{-1}$ [40] represent the scale factor and Hubble parameter today. $a_{\kappa}$ and $H_{\kappa}$ represent the scale factor and Hubble parameter at the pivot scale $\kappa_{*}=0.05\hskip 2.84526pt\text{Mpc}^{-1}$. The subscripts ’eq’, ’end’, and ’reh’ correspond to ’matter-radiation equality’, ’end of inflation’, and ’reheating’, respectively. Note that we solve Eq. (3.2) by using $T_{\text{eq}}=0.8$ eV, and the end of preheating is evaluated when the amplitude of $\phi$, which is $\tilde{\phi}$, equals $M$. On the other hand, we used $\frac{a_{\text{eq}}H_{\text{eq}}}{a_{0}H_{0}}\simeq\sqrt{1+z_{\text{eq}}}\simeq 3400$ by assuming the matter domination near equality.
Finally, by using (3.2) with e-fold $N=75.5$, we obtain an upper bound for the reheating temperature as $T_{\text{reh}}\sim 10^{4}$ GeV. This value is five orders of magnitude lower than the previous prediction [9]. The reheating temperature in our calculation is evaluated at the early radiation-dominated regime.
One should note that even a small change in $n_{s}$ can lead to a substantial variation in the number of e-folds and, consequently, a significant change of the reheating temperature. In this work, we argue that the reheating temperature is directly connected to the preheating stage. As a results, the following sections explain how both the preheating stage and the reheating mechanism are adjusted to accommodate the new, lower reheating temperature favored by ACT results.
@@ 4 The preheating
When $\tilde{\phi}\lesssim M_{P}$, the inflaton oscillates, marking the end of inflation. The particle produced by this mode is non-perturbative. If we assume $\chi$ is dominantly produced during this mode, which is presented by the action:
$S_{\chi}=\int d^{4}x\sqrt{-g_{J}}\left[-\frac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\frac{1}{2}m_{\chi}^{2}\chi^{2}-\frac{1}{2}\xi R_{J}\chi^{2}\right],$ (4.1)
we can evaluate its particle production. Firstly, we write the analytical solution of $\phi$ as [35]
@@
$\ddot{\chi}_{k}+3H\dot{\chi}_{k}-\sqrt{\frac{2}{3}}\frac{\dot{\phi}\dot{\chi}_{k}}{M_{p}}+\left[\frac{k^{2}_{\chi}}{a^{2}}+e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}m_{\chi}^{2}+3\xi M^{2}\left(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}\right)\right]\chi_{k}=0.$ (4.3)
@@
on calculating (4.3). In Fig. 2, we show that the resonance production of $\chi_{k}$ is highly sensitive to both the non-minimal coupling $\xi$ and the initial condition $\chi_{k}(0)$. If these conditions are not satisfied, Starobinsky inflation faces two possible results: (1) inconsistency with the observed particle abundance today (especially if $\chi$ is related to dark matter) or (2) prolonged preheating to generate the required number of particles.
@@
At the early stage of preheating, the inflaton perturbative decay and the scattering process are possible. The single perturbative decay, in principle, is more favored than scattering, as it needs a large density of incoherent inflaton particles to scatter. In contrast, at the start of preheating, the inflaton is mostly a coherent oscillating field. The perturbative decay itself is slow to be produced at the early stage of preheating. Additionally, if the decay occurs during inflation, it may spoil the slow-roll condition or leave observational relics. On the other hand, the favored $\chi\sim 10^{-2.5}M_{p}$ could be obtained from the fluctuation of $\delta\chi$. But the fluctuation can only provide as large as $H/2\pi\sim 10^{-6}M_{p}$, in which about three orders smaller from the desired value. Lastly, the only possible way came from the fact that $\chi$ should be the spectator field during inflation. On the other hand, as $\chi$ is small-coupled with scalaron, it can be assumed to be saved as a spectator field. We can also conclude that for the successful preheating of the Starobinsky inflation, the appearance of the spectator field is necessary. At the early stage of preheating, in order to fulfill the non-relativistic conditions and the efficient preheating, $k_{\chi}$ should be extremely small. Thus, it gives $k_{\chi}^{2}\ll\xi M^{2}$.
Before proceeding, we clarify why the analytical time evolution of $\phi$ is used, even though it could in principle be numerically coupled with $\chi$ to obtain exact results. First, in the pure Starobinsky model, the scalaron $\phi$ is only weakly coupled to other fields during the initial stage of preheating. In the Einstein frame, the Lagrangian of the $\chi$ field is heavily suppressed by the factor $e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}}$, which indicates that $\phi$ and $\chi$ are weakly coupled. This justifies treating $\chi$ as a spectator field during inflation, so that the analytical form of Eq. (4.2) remains valid until $\phi$ becomes small. Also, we restrict our attention to the early phase, where $\phi$ is still not too small after the first crossing. In this regime, the analytical estimation of the time evolution of $\phi$ remains reliable. Second, solving the coupled equations of motion poses a significant challenge, as it involves complicated calculations, including integro-differential equations in $\chi^{2}$. Third, since our paper primarily focuses on the reheating mechanism in light of the new ACT results, we approximate the preheating by considering how efficiently it can drain the inflaton field energy density. A more detailed analysis of the preheating stage may be presented in the future work.
@@
The approximation adopted in Eq. (4.5) is likely oversimplified. As illustrated in Fig. 3, the analytical solution exhibits a significant deviation from the full equation of motion given in Eq. (4.3). In that figure, we present a comparison of $\chi_{k}^{2}$ rather than $\chi_{k}$ itself, since it directly reflects the produced energy density $\delta\rho_{\text{cross}}\propto|\chi_{k}|^{2}$. Figure 3 shows that the numerically obtained energy density of the $\chi$ field is approximately one half of the analytical estimate. We will use this results to estimate particle production during the crossing from the analytical calculations.
Figure 3: The comparison of analytical result (solid-red) and numerical result (solid-black) in $\chi_{k}^{2}$. In the figure, $\chi_{k}^{2}$ of the analytical result is shown twice of the results compared by the numerical ones. In this plot, we used $\xi=10$, $k_{\chi}=10^{-8}M_{p}$, $m_{\chi}=10^{-8}M_{p}$, $\chi_{k}(0)=10^{2.5}M_{p}$ and $\dot{\chi}_{k}=0$. The unit in the $\chi_{k}^{2}$-axis is in order of $M_{p}^{2}$, meanwhile the $t$-axis is in $M_{p}^{-1}$.
To estimate the analytical assumption, we need to focus on the particle production during the first near-zero crossing ($\sin(Mt)\simeq Mt$). Firstly, we can approximate Eq. (4.5) as
@@
$\begin{split}&p^{2}=\left(k_{\chi}^{2}+m_{\chi}^{2}\right)\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{-2/3},\\ &q=\left[\sqrt{\frac{2}{3}}\left(m_{\chi}^{2}+3\xi M^{2}\right)\frac{\tilde{\phi}}{M_{P}}M\right]^{1/3}t.\end{split}$ (4.7)
@@
For $p^{2}\ll q$, we can estimate the particle production of $\chi$ by the first crossing of $\phi$ as [6, 52]
@@
In this calculation, for the crossing to achieve the same energy as the inflaton ($\tfrac{1}{2}M^{2}\tilde{\phi}^{2}\sim 10^{-11}M_{p}^{4}$), we require $\xi\gg\mathcal{O}(1)$, which is strongly disfavored due to naturalness issues. In addition, Fig. 3 shows that the actual value of $\delta\rho_{\text{cross}}$ obtained numerically is about half of the analytical results. These results are largely due to our assumption that $a=1$.
@@
Finally, we obtain the famous Mathieu equation as
@@
where $\theta_{\mu}$ is a phase. It is shown in the Mathieu instability band of Fig. 4 that certain parameters could potentially fall in the instability region in which the particle is largely produced. However, this semi-analytical result of the instability band could not show the effect of the initial conditions. Moreover, the parameters used in the numerical calculation can be mapped onto Fig. 4, with some degree of deviation.